### v1

@@ 5 GFT-Optimal Simple Mechanisms under Independence
In the previous section we have shown that for unrestricted joint distributions, it is impossible to learn an approximately GFT-optimal simple mechanism. Given this impossibility result for arbitrary joint distributions, we move to focus on product distributions, where the valuations of the agents that are sampled independently. Thus, we assume that the value of seller $s$ is sampled from $V_{s}$, and the value of buyer $b_{i}\in\{b_{1},b_{2}\}$ is sampled from $V_{i}$, so the triplet $(v_{s},v_{1},v_{2})$ is sampled from the product distribution $\textit{{V}}=V_{s}\times V_{1}\times V_{2}$. In this section we establish that under independence, the GFT of any simple mechanism can also obtained by a simple mechanism characterized by a compatible pair of associated functions which are monotone, and with a specific structure (being “tight”). These properties of the pair of functions will enable learning a simple mechanism with GFT that is close to being GFT-optimal, as described in the following sections.
@@
[…]
@@ 6 Computing a GFT-optimal Simple Mechanism under Independence
In this section we consider the problem of finding a GFT-optimal mechanism for a given finite product distribution $\textit{{V}}=V_{s}\times V_{1}\times V_{2}$. First, to aid our description and computation of the algorithm, we construct an ordered set $S=\{s_{1}<...<s_{m}\}$ of size $m$, which is a union of the finite supports of the distributions $V_{s},V_{1},V_{2}$, so that the realized value of every agent is always in $S$. Clearly, it is sufficient to define the value of the functions $f_{1}^{*}$ and $f_{2}^{*}$ on on values in $S$. Moreover, we claim that to maximize the GFT, it is sufficient to consider only functions that output values in $S$. This is so because the mechanisms we consider break ties in favor of trade and in favor of the higher value. Therefore, rounding a price that is not in $S$ to the closest value in $S$ would result in the exact same trade in all cases. This fact also allows us to assume that the (unrestricted) best prices $p_{1},p_{2}$ are in $S$.
@@
Since at this step we only need to find the functions $f_{1},f_{2}$ from $(p_{1},p_{2})$ on-wards and we know that they pass through that point, this condition is met at $(p_{1},p_{2})$ and so if the matrix is filled in correctly we will be able to use it to construct the functions $f_{1},f_{2}$ from $(p_{1},p_{2})$ up to $(s_{m},s_{m})$. Figure 10 illustrates the two parts of the functions that we compute in the algorithm (up to $(p_{1},p_{2})$, and from there onward). The definition of $G[i,j]$ requires considering all compatible pairs of functions that begin at $s_{i},s_{j}$, and there are exponentially many such compatible pairs of functions. However, in the next subsection we will show how the matrix can nevertheless be filled in polynomial time.
@@ 7 Learning to Maximize GFT under Independence
Definition 7.1 (Def 14.6 from [MU17]).
@@ B.1 Upper Semi-Continuity of GFT
Theorem 5 (Continuity from above, Theorem 2.1 in [Bil95]).
Let $(\Omega,\mathcal{F},P)$ be a probability space, and let $A_{1}\supseteq A_{2}\supseteq...$ be a sequence of subsets of $\Omega$ ($\forall i\ A_{i}\subseteq\Omega$). Define:
@@
Let $(\Omega,\mathcal{F},P)$ be a probability space, and let $A_{1}\supseteq A_{2}\supseteq...$ be a sequence of subsets of $\Omega$ ($\forall i\ A_{i}\subseteq\Omega$). Then $\lim_{n\rightarrow\infty}P(A_{n})-P(A)=0$. Additionally, if $\Omega\subseteq\mathbb{R}^{k}$ then:
@@ B.2 Restricted Best Price
By Lemma B.5 we get that $\operatorname*{arg\,max}\limits_{p\in\mathbb{R}}\textup{GFT}(p,V_{s}\times V_{b})$ is well defined. Now, towards a contradiction, assume that there is no $\max\left\{\operatorname*{arg\,max}\limits_{p\in\mathbb{R}}\textup{GFT}(p,V_{s}\times V_{b})\right\}$. Therefore, we have an infinite increasing series of $p_{1}<p_{2}<...$ s.t. $\textup{GFT}(p_{1},V_{s}\times V_{b})=\textup{GFT}(p_{2},V_{s}\times V_{b})=...$. Again using the fact that GFT is upper semi-continuous at $p$, we know that the supremum of this sequence is attainable, and that there is a $\max\left\{\operatorname*{arg\,max}\limits_{p\in\mathbb{R}}\textup{GFT}(p,V_{s}\times V_{b})\right\}$. ∎
@@ Appendix C Missing Proofs From Section 7
Definition C.1 (Def 14.1 from [MU17]).
@@
In this paper we use the range space where $X=[0,1]$ and $\mathcal{R}$ is the family of all closed intervals $[a,b]\subseteq[0,1]$. It is known that the VC dimension of this range space is $2$ ([MU17], page 364).
@@
Theorem 6 (Thm. 14.15 from [MU17]).
@@
Now let $S$ be the union of the supports of $S_{s},S_{1},S_{2}$ with $\{0,1\}$, and denote its elements by $S=\{0=s_{1}<...<s_{k}=1\}$. Fix $(s_{i-1},s_{i}]$ and let $v_{1}\in(s_{i-1},s_{i}]$:

### v2

@@ 5 GFT-Optimal Simple Mechanisms under Independence
In the previous section we have shown that for unrestricted joint distributions, it is impossible to learn an approximately GFT-optimal simple mechanism. Given this impossibility result for arbitrary joint distributions, we move to focus on product distributions, where the valuations of the agents that are sampled independently. Thus, we assume the value of seller $s$ is sampled from $V_{s}$, and the value of buyer $b_{i}\in\{b_{1},b_{2}\}$ is sampled from $V_{i}$, so the triplet $(v_{s},v_{1},v_{2})$ is sampled from the product distribution $\textit{{V}}=V_{s}\times V_{1}\times V_{2}$. In this section we establish that under independence, the GFT of any simple mechanism can also obtained by a simple mechanism characterized by a compatible pair of associated functions which are monotone, and with a specific structure (being “tight”). These properties of the pair of functions will enable learning a simple mechanism with GFT that is close to being GFT-optimal, as described in the following sections.
@@
$r_{b}^{f_{-b}}(v_{-b})=\begin{cases}\sup\{v_{b}|v_{-b}\geq f_{-b}(v_{b})\}&\text{if }\exists v_{b}:v_{-b}\geq f_{-b}(v_{b})\\ 0&\text{otherwise}\end{cases}$
@@ 6 Computing a GFT-optimal Simple Mechanism under Independence
In this section we consider the problem of finding a GFT-optimal mechanism for a given finite product distribution $\textit{{V}}=V_{s}\times V_{1}\times V_{2}$. First, to aid our description and computation of the algorithm, we construct an ordered set $S=\{s_{1}<\ldots<s_{m}\}$ of size $m$, which is a union of the finite supports of the distributions $V_{s},V_{1},V_{2}$, so that the realized value of every agent is always in $S$. Clearly, it is sufficient to define the value of the functions $f_{1}^{*}$ and $f_{2}^{*}$ on on values in $S$. Moreover, we claim that to maximize the GFT, it is sufficient to consider only functions that output values in $S$. This is so because the mechanisms we consider break ties in favor of trade and in favor of the higher value. Therefore, rounding a price that is not in $S$ to the closest value in $S$ would result in the exact same trade in all cases. This fact also allows us to assume that the (unrestricted) best prices $p_{1},p_{2}$ are in $S$.
@@
Since at this step we only need to find the functions $f_{1},f_{2}$ from $(p_{1},p_{2})$ on-wards and we know that they pass through that point, this condition is met at $(p_{1},p_{2})$ and so if the matrix is filled in correctly we will be able to use it to construct the functions $f_{1},f_{2}$ from $(p_{1},p_{2})$ up to $(s_{m},s_{m})$. Figure 10 illustrates the two parts of the functions that we compute in the algorithm (up to $(p_{1},p_{2})$, and from there onward). The definition of $G[i,j]$ requires considering all compatible pairs of functions that begin at $s_{i},s_{j}$, and there are exponentially many such compatible pairs of functions. However, in the next subsection we show how the matrix can nevertheless be filled in polynomial time.
@@ 7 Learning to Maximize GFT under Independence
Definition 7.1 (Def 14.6 from [mitzenmacher2017probability]).
@@ B.1 Upper Semi-Continuity of GFT
Theorem 5 (Continuity from above, Theorem 2.1 in [billingsley1995probability]).
Let $(\Omega,\mathcal{F},P)$ be a probability space, and let $A_{1}\supseteq A_{2}\supseteq\ldots$ be a sequence of subsets of $\Omega$ ($\forall i\ A_{i}\subseteq\Omega$). Define:
@@
Let $(\Omega,\mathcal{F},P)$ be a probability space, and let $A_{1}\supseteq A_{2}\supseteq\ldots$ be a sequence of subsets of $\Omega$ ($\forall i\ A_{i}\subseteq\Omega$). Then $\lim_{n\rightarrow\infty}P(A_{n})-P(A)=0$. Additionally, if $\Omega\subseteq\mathbb{R}^{k}$ then:
@@ B.2 Restricted Best Price
By Lemma B.5 we get that $\operatorname*{arg\,max}\limits_{p\in\mathbb{R}}\textup{GFT}(p,V_{s}\times V_{b})$ is well defined. Now, towards a contradiction, assume that there is no $\max\left\{\operatorname*{arg\,max}\limits_{p\in\mathbb{R}}\textup{GFT}(p,V_{s}\times V_{b})\right\}$. Therefore, we have an infinite increasing series of $p_{1}<p_{2}<\ldots$ s.t. $\textup{GFT}(p_{1},V_{s}\times V_{b})=\textup{GFT}(p_{2},V_{s}\times V_{b})=\ldots$. Again using the fact that GFT is upper semi-continuous at $p$, we know that the supremum of this sequence is attainable, and that there is a $\max\left\{\operatorname*{arg\,max}\limits_{p\in\mathbb{R}}\textup{GFT}(p,V_{s}\times V_{b})\right\}$. ∎
@@ Appendix C Missing Proofs From Section 7
Definition C.1 (Def 14.1 from [mitzenmacher2017probability]).
@@
In this paper we use the range space where $X=[0,1]$ and $\mathcal{R}$ is the family of all closed intervals $[a,b]\subseteq[0,1]$. It is known that the VC dimension of this range space is $2$ ([mitzenmacher2017probability], page 364).
@@
Theorem 6 (Thm. 14.15 from [mitzenmacher2017probability]).
@@
Now let $S$ be the union of the supports of $S_{s},S_{1},S_{2}$ with $\{0,1\}$, and denote its elements by $S=\{0=s_{1}<\ldots<s_{k}=1\}$. Fix $(s_{i-1},s_{i}]$ and let $v_{1}\in(s_{i-1},s_{i}]$: